A woman jogs for 3 km at 9 km/h and then walks for 2 km at 5 km/h. What is her average speed for the entire trip?
Answer #1

Time, t1 = 3/9 = 1/3 = 0.33 hrs

Time, t2 = 2/5 = 0.4 hrs

Average speed, v = (3 + 2)/(0.33 + 0.4) = 6.85 km/hr

Comment in case any doubt please rate my answer. ...